X Company is considering buying a part in 2020 that it has been making for the past several years. A company has offered to supply this part for $15.93 per unit. Budgeted production in 2020 is 54,000 units, and budgeted per-unit production costs are: Materials $5.90 Direct labor [all variable] 5.70 Total overhead 4.40 Total $16.00 $102,600 of X Company's total overhead costs are fixed; $73,872 of the $102,600 are unavoidable even if it buys the part. Also, if X Company buys the part, it can rent the equipment that was used to make the part to another company for $80,000.

If X Company buys the part instead of making it, it will save

At what production level would X company be indifferent between making and buying the part?
Answer #1

a) Differential analysis

Make Buy
Direct material 318600
Direct labor 307800
Variable overhead 135000
Fixed overhead 28728
Opportunity Cost 80000
Purchase cost 860220
Total relevant cost 870128 860220

it will save = 870128-860220 = 9908

b) Indifferent point

Manufacturing cost = Purchase cost

14.10X+108728 = 15.93

-1.83X = 108728

X(Unit) = 59414 Units